# ContextGPT: Infusing LLMs Knowledge into Neuro-Symbolic Activity Recognition Models

### v1

@@ Abstract.
Context-aware Human Activity Recognition (HAR) is a hot research area in mobile computing, and the most effective solutions in the literature are based on supervised deep learning models. However, the actual deployment of these systems is limited by the scarcity of labeled data that is required for training. Neuro-Symbolic AI (NeSy) provides an interesting research direction to mitigate this issue, by infusing common-sense knowledge about human activities and the contexts in which they can be performed into HAR deep learning classifiers. Existing NeSy methods for context-aware HAR rely on knowledge encoded in logic-based models (e.g., ontologies) whose design, implementation, and maintenance to capture new activities and contexts require significant human engineering efforts, technical knowledge, and domain expertise. Recent works show that pre-trained Large Language Models (LLMs) effectively encode common-sense knowledge about human activities. In this work, we propose ContextGPT: a novel prompt engineering approach to retrieve from LLMs common-sense knowledge about the relationship between human activities and the context in which they are performed. Unlike ontologies, ContextGPT requires limited human effort and expertise. An extensive evaluation carried out on two public datasets shows how a NeSy model obtained by infusing common-sense knowledge from ContextGPT is effective in data scarcity scenarios, leading to similar (and sometimes better) recognition rates than logic-based approaches with a fraction of the effort.
@@
human activity recognition, context-awareness, large language models, neuro-symbolic, knowledge infusion
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learningccs: Human-centered computing Ubiquitous and mobile devicesccs: Information systems Information extraction
1. Introduction
The analysis of sensor data gathered from mobile and wearable devices for Human Activity Recognition (HAR) has been extensively researched (Chen et al., 2021; Gu et al., 2021). This is attributed to its high potential for applications across various domains such as well-being, healthcare, and sports, fueled by the widespread adoption of these devices.
@@ 1. Introduction
Although the majority of the studies in this area focused on inertial sensors only, context-aware HAR approaches also take into account the user’s surrounding context as derived from the devices, user profiles or online services (e.g., semantic location, temporal information, etc.) to further improve the recognition rate and, at the same time, extend the set of recognizable activities (Bettini et al., 2020).
@@
In the last few years, a vast amount of solutions based on deep learning classifiers have been proposed (Wang et al., 2019). However, most of those approaches rely on supervised learning and require large labeled datasets to be trained. The need to acquire reliable labels from a large number of users and for an extended set of activities is currently a major obstacle to the deployment of effective sensor-based HAR systems.
@@
Among the several solutions proposed in the general machine learning community to tackle the labeled data scarcity problem, Neuro-Symbolic (NeSy) approaches represent a promising research direction (Sarker et al., 2021). NeSy methods aim at combining data-driven and knowledge-based approaches to reduce the amount of labeled data needed to effectively train the model and, at the same time, to improve its interpretability. In NeSy approaches, well-known domain constraints are infused into the model, thus avoiding learning them from data and simplifying the overall learning process.
@@
Existing NeSy approaches proposed for context-aware HAR retrieve common-sense knowledge from logic-based models (e.g., ontologies) manually designed and implemented by human experts (Arrotta et al., 2022). Such knowledge models encode the relationship between an activity and the context in which it can be performed. For instance, they encode that biking is not a plausible activity if the user’s semantic location is highway or museum. However, manually building a comprehensive knowledge model that captures all possible context situations is challenging, and, even relying on a domain expert, it does not guarantee that all the possible situations in which an activity can be performed are captured. Moreover, logic-based models are not scalable since including new activities, new context conditions, or new constraints requires further manual work and expertise.
@@
This paper explores the idea of infusing common-sense knowledge into NeSy HAR models from Large Language Models (LLMs) instead of ontologies. Indeed, LLMs implicitly encode knowledge spanning a wide range of domains and recent results suggest that they can be efficiently exploited to retrieve common-sense knowledge about human activities (Takeda et al., 2023; Graule and Isler, 2023; Zhou et al., 2023; Kaneko and Inoue, 2023; Xia et al., 2023; Leng et al., 2023).
@@
2. Related Work
2.1. Data scarcity in HAR
Most of the solutions proposed in the literature for sensor-based HAR on mobile/wearable devices rely on supervised Deep Learning (DL) methods (Wang et al., 2019; Chen et al., 2021). Even though the majority of these works only focused on inertial sensors, several studies highlight how also including high-level context information can significantly improve the recognition rate(Bettini et al., 2020; Saguna et al., 2013).
@@ 2.1. Data scarcity in HAR
Due to the unrealistic amount of training data required to train supervised models, several research groups are proposing solutions to leverage small amounts of labeled samples. Proposed methods to mitigate labeled data scarcity are based on transfer learning (Sanabria et al., 2021; Soleimani and Nazerfard, 2021), self-supervised learning (Haresamudram et al., 2022; Jain et al., 2022; Hiremath et al., 2022), and semi-supervised learning approaches (Abdallah et al., 2018). We believe that Neuro-Symbolic AI (NeSy) could be coupled with such approaches to further mitigate data scarcity when fine-tuning pre-trained HAR models with limited labeled data.
2.2. Neuro-Symbolic HAR
@@
While several NeSy methods have been proposed in the computer vision and NLP domains (Dash et al., 2022), only a few approaches have been proposed for HAR.
@@
Existing NeSy methods for context-aware HAR retrieve common-sense knowledge from logic-based models (e.g., ontologies). To the best of our knowledge, three main strategies have been proposed so far to combine extracted knowledge with deep learning models: a) using knowledge to refine the deep model’s output (Bettini et al., 2020), b) including retrieved knowledge as additional features in the latent space (Arrotta et al., 2020), and c) using a loss function that penalizes predictions violating domain constraints (Arrotta et al., 2023a; Xing et al., 2020). However, designing and implementing knowledge models require significant human effort, and those models may not capture all the possible situations in which activities can be performed. While there are information retrieval approaches to semi-automatically obtaining common-sense knowledge from public external sources (e.g., images (Riboni and Murtas, 2019), web (Perkowitz et al., 2004), text (Yordanova, 2016)) such methods still faces challenges in creating comprehensive knowledge models.
2.3. Using LLMs for Human Activity Recognition
@@
The adoption of Large Language Models (LLMs) in sensor-based HAR is a recently emerging trend, and we expect a surge of contributions in this area in the near future. For instance, the work in (Takeda et al., 2023) takes advantage of the GPT-2 model to predict sequences of sensor events in smart-home settings. Specifically, since the GPT-2 model is pre-trained to predict the next word in a sentence, it has been fine-tuned to predict the most likely sequence of sensor events that are treated similarly to textual data. A similar approach was also proposed in (Graule and Isler, 2023), leveraging the LLaMA2 language model. The work in (Zhou et al., 2023) uses the CLIP pre-trained model to align sensor data embeddings and text embeddings by means of contrastive learning, with the goal of improving the overall recognition rate.
@@
A few research groups recently proposed solutions based on the well-known ChatGPT tool. In (Kaneko and Inoue, 2023), ChatGPT is prompted with the description of a specific sensor-based HAR task and the current choice for sensor positioning, with the objective of obtaining feedback about where to install new sensors to further improve the recognition rate. The work in (Xia et al., 2023) proposed an unsupervised approach for smart-home settings. The approach consists of two steps: the former asks ChatGPT to provide a description of the target activities, and the latter asks ChatGPT to infer the performed activity given a temporal sequence of events describing interaction with household items. Finally, based on the recent advances in generating synthetic inertial sensor data from textual descriptions (Guo et al., 2022), the work in (Leng et al., 2023) leverages ChatGPT to generate textual descriptions of activities, that are then used to generate virtual inertial sensors data.
To the best of our knowledge, we are the first to leverage LLMs to mitigate data scarcity in context-aware HAR based on mobile/wearable devices. In particular, we query LLMs to determine which activities are compatible with the current user’s context, with the goal of infusing this knowledge into a neuro-symbolic model.
3. Methodology
3.1. A Neuro-Symbolic Framework for Context-Aware HAR
We consider a mobile and wearable computing scenario in which users perform activities while carrying their devices (e.g., smartphone, smartwatch, tracker).
@@ 3.1. A Neuro-Symbolic Framework for Context-Aware HAR
Figure 1 provides an overview of the Neuro-Symbolic (NeSy) system taking as input the continuous stream of sensor data generated by the user devices, and providing as output the most likely performed activity.
@@
Figure 1. A Neuro-Symbolic AI framework for context-aware HAR gathering knowledge from ContextGPT
@@
As a common practice in this field (Chen et al., 2021), data are partitioned into temporal windows. For each window $w$ (representing $z$ seconds of consecutive raw sensor data), we derive two subsets $w^{R}$ and $w^{C}$: the former includes raw data that we consider appropriate for being directly processed by a data-driven model (e.g., inertial sensors data), while the latter involves raw sensor data that we consider useful for deriving high-level contexts through reasoning and/or abstraction.
@@
The Context Aggregation module is in charge of deriving all the most likely high-level contexts $C^{w}\subset C$ that occur in a window $w$ based on $w^{C}$. Context Aggregation can be implemented using simple rules, available services, and/or context-aware middlewares (Henricksen et al., 2005). For example, raw GPS coordinates can be used to derive the semantic location by querying a dedicated web service (e.g., by using Google Places APIs).
Context-aware HAR could be addressed by relying on machine learning models taking $\langle w^{R},C^{w}\rangle$ only as input. However, a more effective approach is to complement data-driven approaches with common-sense knowledge about the relationships between activities and contexts (Riboni and Bettini, 2011). For example, based on common-sense knowledge, people typically run outdoors in parks, on trails, or along sidewalks (preferably in dry weather) and indoors on a treadmill. Moreover, running requires the user to move at a positive speed. The relationships between activities and the typical conditions in which they can be performed can be used in the HAR process to reduce the amount of labeled data required to learn them.
The ContextGPT module (that is the main contribution of this work) is in charge of reasoning on the contexts in $C^{w}$ to derive the most likely activities that are consistent according to $C^{w}$. ContextGPT relies on a Large Language Model (LLM) and it is described in detail in the rest of this section. The information about context-consistent activities is infused into a NeSy HAR model, which also receives as input raw data and high-level context data (i.e. $\langle w^{R},C^{w}\rangle$). The output of the model is a probability distribution $P=\langle p_{1},\dots,p_{k}\rangle$, where $p_{i}$ is the probability that the user performed the activity $a_{i}$.
3.2. ContextGPT architecture
@@
Figure 2. Overall architecture of ContextGPT
ContextGPT receives as input a temporal window of high-level context data $C^{w}$. First, $C^{w}$ is provided to the Context2Text module to obtain a natural language description of the user’s context. Since LLMs also benefit from a few examples depicting how the required task should be carried out (i.e., the so-called few-shot prompting (Brown et al., 2020)), the Example selection module considers a pool of examples and includes in the prompt those having their context similar to $C^{w}$. Each example depicts a context situation and the activities that are consistent with that situation. Finally, the Prompt Construction module generates the complete prompt that is composed of:
@@
Figure 3 shows an example of a typical prompt generated by ContextGPT.
Figure 3. An example of ContextGPT prompt
@@
3.3. Prompt Engineering
While existing Neuro-Symbolic methods for Context-Aware HAR demand knowledge engineers to manually build logic-based models, ContextGPT requires a Prompt Engineer in charge of designing: i) the system message, ii) the translation of context data into natural language descriptions, and iii) the examples in the pool. Due to the sensitivity of LLMs to the specific wording in the prompt, the Prompt Engineer currently has a non-trivial role for the success (or failure) in obtaining the desired goal (Meyer et al., 2023). However, since these tasks are based on natural language, there is no need for designing comprehensive and complex relationships between activities and contexts using logic formalisms; hence, the required expertise and human effort is significantly reduced. In the following, we describe each component of the prompts of ContextGPT in detail.
3.3.1. System Message
The system message defines the general description of the task the LLM should complete. In our case, the task is determining the activities that are consistent with a given context. Hence, we divided the system message into two parts. The former instructs the LLM about the overall task and the list of possible activities. The latter provides a detailed list of steps the LLM should undertake to complete the task (i.e., Chain-Of-Thought approach (Wei et al., 2022)).
The first step directs the LLM to focus on the input context. The second step requires following an open-world assumption since, in our preliminary experiments, we noticed instances where the model mistakenly excluded activities not explicitly supported by the context description. For instance, consider the following context description: “In the last 4 seconds the user Bob was in an outdoor environment, where he was moving/traveling at speed between 1 and 4 km/h, experiencing rainy weather, not following/close to a public transportation route, and not experiencing elevation changes.”. Without the second step, in our preliminary experiments, the LLM often excluded the activity Moving by Car with the following motivation: “Not consistent as there is no mention of being in a car or any other vehicle besides walking speed movement”. While it is true that being in a car is not explicitly provided in the context description, the Moving by car activity should still be considered as possible. Indeed, it is impractical to monitor all the possible contextual conditions through sensors in mobile/wearable devices. Finally, the last step forces the LLM to provide context-consistent activities in a specific format to simplify the automatic extraction of this information during post-processing.
Figure 4 shows the system message of ContextGPT.
Figure 4. The system message of ContextGPT. The possible activities, in this case, are the ones of the DOMINO (Arrotta et al., 2023b) dataset.
3.3.2. Context2Text

### v2

@@ Abstract
Context-aware Human Activity Recognition (HAR) is a hot research area in mobile computing, and the most effective solutions in the literature are based on supervised deep learning models. However, the actual deployment of these systems is limited by the scarcity of labeled data that is required for training. Neuro-Symbolic AI (NeSy) provides an interesting research direction to mitigate this issue, by infusing common-sense knowledge about human activities and the contexts in which they can be performed into HAR deep learning classifiers. Existing NeSy methods for context-aware HAR rely on knowledge encoded in logic-based models (e.g., ontologies) whose design, implementation, and maintenance to capture new activities and contexts require significant human engineering efforts, technical knowledge, and domain expertise. Recent works show that pre-trained Large Language Models (LLMs) effectively encode common-sense knowledge about human activities. In this work, we propose ContextGPT: a novel prompt engineering approach to retrieve from LLMs common-sense knowledge about the relationship between human activities and the context in which they are performed. Unlike ontologies, ContextGPT requires limited human effort and expertise, while sharing similar privacy concerns if the reasoning is performed in the cloud. An extensive evaluation using two public datasets shows how a NeSy model obtained by infusing common-sense knowledge from ContextGPT is effective in data scarcity scenarios, leading to similar (and sometimes better) recognition rates than logic-based approaches with a fraction of the effort.
Index Terms:
@@
human activity recognition, context-awareness, large language models
I Introduction
@@
The analysis of sensor data gathered from mobile and wearable devices for Human Activity Recognition (HAR) has been extensively researched [1, 2]. This is attributed to its high potential for applications across various domains such as well-being, healthcare, and sports, fueled by the widespread adoption of these devices. Although the majority of the studies in this area focused on inertial sensors only, context-aware HAR approaches also take into account the user’s surrounding context as derived from the devices, user profiles or online services (e.g., semantic location, temporal information, etc.) to further improve the recognition rate and, at the same time, extend the set of recognizable activities [3].
@@
In the last few years, a vast amount of solutions based on deep learning classifiers have been proposed [4]. However, most of those approaches rely on supervised learning and require large labeled datasets to be trained. The need to acquire reliable labels from a large number of users and for an extended set of activities is currently a major obstacle to the deployment of effective sensor-based HAR systems.
@@
Among the several solutions proposed in the general machine learning community to tackle the labeled data scarcity problem, Neuro-Symbolic (NeSy) approaches represent a promising research direction [5]. NeSy methods aim at combining data-driven and knowledge-based approaches to reduce the amount of labeled data needed to effectively train the model and, at the same time, to improve its interpretability. In NeSy approaches, well-known domain constraints are infused into the model, thus avoiding learning them from data and simplifying the overall learning process.
@@
Existing NeSy approaches proposed for context-aware HAR retrieve common-sense knowledge from logic-based models (e.g., ontologies) manually designed and implemented by human experts [6]. Such knowledge models encode the relationship between an activity and the context in which it can be performed. For instance, they encode that biking is not a plausible activity if the user’s semantic location is highway or museum. However, manually building a comprehensive knowledge model that captures all possible context situations is challenging, and, even relying on a domain expert, it does not guarantee that all the possible situations in which an activity can be performed are captured. Moreover, logic-based models are not scalable since including new activities, new context conditions, or new constraints requires further manual work and expertise.
@@
This paper explores the idea of infusing common-sense knowledge into NeSy HAR models from Large Language Models (LLMs) instead of ontologies. Indeed, LLMs implicitly encode knowledge spanning a wide range of domains and recent results suggest that they can be efficiently exploited to retrieve common-sense knowledge about human activities [7, 8, 9, 10, 11, 12].
@@
II Related Work
II-A Data scarcity in HAR
@@
Due to the unrealistic amount of training data required to train supervised models, several research groups are proposing solutions to leverage small amounts of labeled samples. Proposed methods to mitigate labeled data scarcity are based on transfer learning [13], self-supervised learning [14], and semi-supervised learning approaches [15]. We believe that Neuro-Symbolic AI (NeSy) could be coupled with such approaches to further mitigate data scarcity when fine-tuning pre-trained HAR models with limited labeled data.
II-B Neuro-Symbolic HAR
@@
While several NeSy methods have been proposed in the computer vision and NLP domains [16], only a few approaches have been proposed for HAR.
@@
Existing NeSy methods for context-aware HAR retrieve common-sense knowledge from logic-based models (e.g., ontologies). To the best of our knowledge, three main strategies have been proposed so far to combine extracted knowledge with deep learning models: a) using knowledge to refine the deep model’s output [3], b) including retrieved knowledge as additional features in the latent space [17], and c) using a loss function that penalizes predictions violating domain constraints [18, 19]. However, designing and implementing knowledge models require significant human effort, and those models may not capture all the possible situations in which activities can be performed. While there are information retrieval approaches to semi-automatically obtaining common-sense knowledge from public external sources (e.g., images [20], web [21], text [22]) such methods still faces challenges in creating comprehensive knowledge models.
II-C Using LLMs for Human Activity Recognition
@@
The adoption of Large Language Models (LLMs) in sensor-based HAR is a recently emerging trend, and we expect a surge of contributions in this area in the near future. For instance, the work in [7] takes advantage of the GPT-2 model to predict sequences of sensor events in smart-home settings. Specifically, since the GPT-2 model is pre-trained to predict the next word in a sentence, it has been fine-tuned to predict the most likely sequence of sensor events that are treated similarly to textual data. A similar approach was also proposed in [8], leveraging the LLaMA2 language model. The work in [9] uses the CLIP pre-trained model to align sensor data embeddings and text embeddings by means of contrastive learning, with the goal of improving the overall recognition rate.
@@
A few research groups recently proposed solutions based on the well-known ChatGPT tool. In [10], ChatGPT is prompted with the description of a specific sensor-based HAR task and the current choice for sensor positioning, with the objective of obtaining feedback about where to install new sensors to further improve the recognition rate. The work in [11] proposed an unsupervised approach for smart-home settings. The approach consists of two steps: the former asks ChatGPT to provide a description of the target activities, and the latter asks ChatGPT to infer the performed activity given a temporal sequence of events describing interaction with household items. Finally, based on the recent advances in generating synthetic inertial sensor data from textual descriptions [23], the work in [12] leverages ChatGPT to generate textual descriptions of activities, that are then used to generate virtual inertial sensors data.
To the best of our knowledge, we are the first leveraging LLMs and NeSy models to mitigate data scarcity in context-aware HAR.
III Methodology
III-A A Neuro-Symbolic Framework for Context-Aware HAR
@@
We consider a mobile and wearable computing scenario in which users perform activities while carrying their devices (e.g., smartphone, smartwatch, tracker). Figure 1 provides an overview of the Neuro-Symbolic (NeSy) framework taking as input the continuous stream of sensor data generated by the user devices, and providing as output the most likely performed activity.
@@
Figure 1: A Neuro-Symbolic AI framework for context-aware HAR gathering knowledge from ContextGPT
@@
Data are partitioned in fixed-size temporal windows, and for each window $w$ we derive two subsets $w^{R}$ and $w^{C}$: the former includes raw data that we consider appropriate for being directly processed by a data-driven model (e.g., inertial sensors data), while the latter involves raw sensor data that we consider useful for deriving high-level contexts through reasoning and/or abstraction.
@@
The Context Aggregation module is in charge of deriving all the most likely high-level contexts $C^{w}\subset C$ that occur in a window $w$ based on $w^{C}$. Context Aggregation can be implemented using simple rules, available services, and/or context-aware middlewares. For example, raw GPS coordinates can be used to derive the semantic location by querying a dedicated web service (e.g., by using Google Places APIs).
Context-aware HAR could be tackled with machine learning models taking $\langle w^{R},C^{w}\rangle$ only as input. However, a more effective approach is to complement data-driven approaches with common-sense knowledge about the relationships between activities and contexts [24]. For example, people typically run outdoors in parks, on trails, or along sidewalks (preferably in dry weather) and indoors on a treadmill. Moreover, running requires the user to move at a positive speed. The relationships between activities and the typical conditions in which they can be performed can be used in the HAR process to reduce the amount of labeled data required to learn them.
The ContextGPT module is in charge of reasoning on the contexts in $C^{w}$ to derive the most likely activities that are consistent according to $C^{w}$. ContextGPT relies on a Large Language Model (LLM) and it is described in detail in the rest of this section. The information about context-consistent activities is infused into a NeSy HAR model, which also receives as input raw data and high-level context data (i.e. $\langle w^{R},C^{w}\rangle$). The output of the model is a probability distribution $P=\langle p_{1},\dots,p_{k}\rangle$, where $p_{i}$ is the probability that the user performed the activity $a_{i}$.
III-B ContextGPT architecture
@@
Figure 2: Overall architecture of ContextGPT
ContextGPT receives as input a temporal window of high-level context data $C^{w}$. First, $C^{w}$ is provided to the Context2Text module to obtain a natural language description of the user’s context. Since LLMs also benefit from a few examples depicting how the required task should be carried out (i.e., the so-called few-shot prompting [25]), the Example selection module considers a pool of examples and includes in the prompt those having their context similar to $C^{w}$. Each example depicts a context situation and the activities that are consistent with that situation. Finally, the Prompt Construction module generates the complete prompt that is composed of:
@@
III-C Prompt Engineering
While existing Neuro-Symbolic methods for Context-Aware HAR demand knowledge engineers to manually build logic-based models, ContextGPT requires a Prompt Engineer in charge of designing: i) the system message, ii) the translation of context data into natural language descriptions, and iii) the examples in the pool. Due to the sensitivity of LLMs to the specific wording in the prompt, the Prompt Engineer currently has a non-trivial role for the success (or failure) in obtaining the desired goal [26]. However, since these tasks are based on natural language, there is no need for designing comprehensive and complex relationships between activities and contexts using logic formalisms; hence, the required expertise and human effort is significantly reduced. In the following, we describe each component of the prompts of ContextGPT in detail.
III-C1 System Message
The system message defines the general description of the task the LLM should complete. In our case, the task is determining the activities that are consistent with a given context. Hence, we divided the system message into two parts. The former instructs the LLM about the overall task and the list of possible activities. The latter provides a detailed list of steps the LLM should undertake to complete the task (i.e., Chain-Of-Thought approach [27]). The first step directs the LLM to focus on the input context. The second step requires following an open-world assumption since, in our preliminary experiments, we noticed instances where the model mistakenly excluded activities not explicitly supported by the context description. For instance, consider the following context description: “In the last 4 seconds the user Bob was in an outdoor environment, where he was moving/traveling at speed between 1 and 4 km/h, experiencing rainy weather, not following/close to a public transportation route, and not experiencing elevation changes.”. Without the second step, in our preliminary experiments, the LLM often excluded the activity Moving by Car with the following motivation: “Not consistent as there is no mention of being in a car or any other vehicle besides walking speed movement”. While it is true that being in a car is not explicitly provided in the context description, the Moving by car activity should still be considered as possible. Indeed, it is impractical to monitor all the possible contextual conditions through sensors in mobile/wearable devices. Finally, the last step forces the LLM to provide context-consistent activities in a specific format to simplify the automatic extraction of this information during post-processing. Figure 3 shows the system message of ContextGPT.
Figure 3: The system message of ContextGPT. The possible activities, in this case, are the ones of the DOMINO [28] dataset.
III-C2 Context2Text